OpenStudy (anonymous):

PROBABILITY & STATISTICS NEED HELP If you separate all the clubs from a deck of cards and pick 3 of those 13 cards, what is the probability that the cards are jack, queen, and king in ascending or descending order?

OpenStudy (anonymous):

I might be oversimplifying, but I think it's just \[\frac{1}{13}\cdot\frac{1}{12}\cdot\frac{1}{11}\] and you double that because you want both ascending and descending sequences. Something tells me there may be more to the problem involving counting permutations. Maybe someone else can shed some light on this.

OpenStudy (anonymous):

thank you thank you

OpenStudy (anonymous):

incorrectto

OpenStudy (anonymous):

Hmm well here's my reasoning. \[\text{desired probability}=\frac{\text{desired outcomes}}{\text{possible outcomes}}\] Of the possible permutations, only 2 fit the description ( JQK and KQJ ), so the numerator would be 2. The total possible would be any 3 cards picked from the deck of 13, or \[\large _{13}P_{3}=\frac{13!}{(13-3)!}=\frac{13!}{10!}=13\times12\times11\] So you have \[P=\frac{2}{13\times12\times11}\] like I said before. Anyone else agree or am I wrong?

OpenStudy (kirbykirby):

Looks right to me

OpenStudy (anonymous):

i clicked it and the only other answer said it was 2,000 somethin. so i told my teacher saying i answerd correct but told me i was wrong its happend before thank you

OpenStudy (anonymous):

The probability of an event is restricted to be between 0 and 1, there's no way it can be that big.

OpenStudy (anonymous):

but uk what i was trying to get across ik that